# When Domains Collide: An Activity Theory Exploration of Cross-Disciplinary Collaboration

### v1

@@ II Background
Collaboration in software development teams is a long-studied phenomenon [15, 16], but as technology and practices evolve, so does how collaboration is conducted, and the frictions involved in it. Here we focus on the collaboration aspect of how SDE works with DEs, which can be across the team border as clients or, as in more recent situations, embedded within the team itself with a shared outcome and client focus. We categorized the different types of collaborations into four models.
@@ III-A Study 1: AT-Guided Exploratory Interview
Pilot interviews. Before conducting the main interviews, we piloted six interviews to collect feedback on the interview design within an organization at Microsoft, involving participants with experience in CDSD collaborations (3 SDEs, 3 DEs). We collected feedback and refined the interview questions accordingly. For example, pilot participants suggested adding follow-up questions to understand the reasons behind expectations.
@@ III-B Study 2: Confirmatory Survey
We also asked participants to rate their agreement on their perceptions of their recent CDSD experiences: “My most recent experience with CDSD in software development was satisfying,”. This question was also inspired by the Linux and Apache survey in understanding contributors’ experience of contribution [51, 52]
@@ IV-B Emerging Expectation in CDSD
(E2) Improving Code Qualities. (Division of Labor) DEs often expect SDEs to improve, refactor, and polish code, “I expect SDEs to speed up and professionalize my code” [P_SDE214]. “They are good at code” [DE10]. SDEs do not always agree with such task distributions “The DEs will need to spend a lot of their time learning…how to be engineers” [SDE3].
@@ V-A Emerging Frictions in CDSD
Frictions from Ambiguous Requirements. SDEs often struggle with ambiguous and evolving requirements. over 70% of SDE participants reported often/occasionally facing this friction. “The project objectives at the start usually do not match the project objectives at the end”[P_SDE114] “Occasionally DEs did not fully understand the proposed solutions before beginning development”[PSDE189]. Ultimately, these ambiguous requests weaken trust and negatively impact collaborations “When the code you’ve just written could be scrapped, and you have to rewrite it in a couple of weeks because they decide: ‘that’s not exactly what we want; we want it to work a little differently” [SDE10].
@@ VI Discussion and Remarks
Implications for Future Research. As shown in Figure 2, Community is heavily implicated with strong paths to Rules and Division of Labor. Future research could: (1) Reimagine hybrid collaboration models that integrate AT with other frameworks such as Distributed Cognition [63], which aims to understand how thinking and knowledge are distributed across people, environments, and tools. Such integration could inform decision-revision processes or adaptive testing workflows that accommodate rapid experimentation while preserving engineering integrity. For example, in CDSD, shared artifacts like decision logs or experiment dashboards could be studied both as mediating tools in AT (surfacing contradictions) and as cognitive scaffolds in DC (distributing decision-making across SDEs and DEs). (2) Investigate how mutual onboarding practices can be designed and evaluated through the lens of Communities of Practice [64], which views learning as a social process rooted in shared practice, to enhance knowledge sharing and accelerate productivity.

### v2

# Zixuan Feng1, Thomas Zimmermann2, Lorenzo Pisani3, Christopher Gooley3, Jeremiah Wander3, Anita Sarma1

@@ II Background
Collaboration in software development teams is a long-studied phenomenon [15, 16], but as technology and practices evolve, so does collaboration and its associated frictions. We focus on the collaboration aspect of how SDE works with DEs, which can be across the team border as clients or, as in more recent situations, embedded within the team itself with a shared outcome and client focus. We categorized the different types of collaborations into four models.
@@ III-A Study 1: AT-Guided Exploratory Interview
Pilot interviews. Before conducting the main interviews, we piloted six interviews to collect feedback on the interview design within an organization at Microsoft, involving participants with experience in CDSD (3 SDEs, 3 DEs). We collected feedback and refined the interview questions accordingly. For example, pilot participants suggested adding follow-up questions to understand the reasons behind expectations.
@@ III-B Study 2: Confirmatory Survey
We also asked participants to rate their agreement with the statement: “My most recent experience with CDSD in software development was satisfying.” This question was inspired by the Linux and Apache surveys on contributors’ experiences [51, 52].
@@ IV-B Emerging Expectation in CDSD
(E2) Improving Code Qualities. (Division of Labor) DEs often expect SDEs to improve, refactor, and polish code, “I expect SDEs to speed up and professionalize my code” [P_SDE214]. “They are good at code” [DE10]. SDEs disagree with such task distributions “The DEs will need to spend a lot of their time learning…how to be engineers” [SDE3].
@@ V-A Emerging Frictions in CDSD
Frictions from Ambiguous Requirements. SDEs often struggle with ambiguous and evolving requirements. over 70% of SDE participants reported often/occasionally facing this friction. “The project objectives at the start usually do not match the project objectives at the end”[P_SDE114] “Occasionally DEs did not fully understand the proposed solutions before beginning development”[P_SDE189]. Ultimately, these ambiguous requests weaken trust and negatively impact collaborations “When the code you’ve just written could be scrapped, and you have to rewrite it in a couple of weeks because they decide: ‘that’s not exactly what we want; we want it to work a little differently” [SDE10].
@@ VI Discussion and Remarks
Implications for Future Research. As shown in Figure 2, Community is heavily implicated with strong paths to Rules and Division of Labor. Future research could: (1) Reimagine hybrid collaboration models that integrate AT with other frameworks such as Distributed Cognition (DC) [63], which aims to understand how thinking and knowledge are distributed across people, environments, and tools. Such integration could inform decision-revision processes or adaptive testing workflows that accommodate rapid experimentation while preserving engineering integrity. For example, in CDSD, shared artifacts like decision logs or experiment dashboards could be studied both as mediating tools in AT (surfacing contradictions) and as cognitive scaffolds in DC (distributing decision-making across SDEs and DEs). (2) Investigate how mutual onboarding practices can be designed and evaluated through the lens of Communities of Practice [64], which views learning as a social process rooted in shared practice, to enhance knowledge sharing and accelerate productivity.